### v2

# Topic-FlipRAG: Topic-Orientated Adversarial Opinion Manipulation Attacks to Retrieval-Augmented Generation Models

Abstract.
Retrieval-Augmented Generation (RAG) systems based on Large Language Models (LLMs) have become essential for tasks such as question answering and content generation. However, their increasing impact on public opinion and information dissemination has made them a critical focus for security research due to inherent vulnerabilities. Previous studies have predominantly addressed attacks targeting factual or single-query manipulations. In this paper, we address a more practical scenario: topic-oriented adversarial opinion manipulation attacks on RAG models, where LLMs are required to reason and synthesize multiple perspectives, rendering them particularly susceptible to systematic knowledge poisoning. Specifically, we propose Topic-FlipRAG, a two-stage manipulation attack pipeline that strategically crafts adversarial perturbations to influence opinions across related queries. This approach combines traditional adversarial ranking attack techniques and leverages the extensive internal relevant knowledge and reasoning capabilities of LLMs to execute semantic-level perturbations. Experiments show that the proposed attacks effectively shift the opinion of the model’s outputs on specific topics, significantly impacting user information perception. Current mitigation methods cannot effectively defend against such attacks, highlighting the necessity for enhanced safeguards for RAG systems, and offering crucial insights for LLM security research.
Corresponding author: laujames2017@whu.edu.cn
1. Introduction
Figure 1. An concise overview of topic-oriented adversarial opinion manipulation attacks on RAG systems.
Retrieval-Augmented Generation (RAG) systems, built on large language models (LLMs), have advanced significantly and been widely applied in tasks like question answering and content generation (Gao et al., 2023; Zhao et al., 2024). These systems integrate retrieval mechanisms with generative models, accessing diverse sources, e.g., Wikipedia, Reddit, and news articles, to deliver updated knowledge. While research has focused on the development of RAG frameworks (Asai et al., 2023; Shi et al., 2025; Xu et al., 2024), the majority of studies emphasize performance and generalization. Consequently, security aspects have received comparatively little attention.
A defining feature of RAG systems is the vast scale of their referenced corpora or knowledge bases. However, these documents might originate from sources beyond service providers’ control and may resist complete content purification. This creates new security risks for RAG-supported LLMs: adversaries can inject meticulously crafted malicious content (Carlini et al., 2024) into retrieval collections, ensuring its prioritized retrieval to influence LLM outputs. The inherent vulnerabilities of RAG models, particularly their persistent impact on public opinion and information dissemination (Chen et al., 2025), have gradually drawn significant attention from the AI safety research community (Zhong et al., 2023; Xue et al., 2024; Zou et al., 2024). Investigating these vulnerabilities is critical for advancing RAG security and ensuring system robustness and credibility.
Early adversarial attacks on RAG systems primarily focused on jailbreak techniques and query-specific perturbations (Chen et al., 2025; Zhong et al., 2023). These methods tamper with knowledge bases to manipulate retrieval rankings, enabling LLMs to provide answers defined by attackers for specific queries. Recent research, such as PoisonedRAG (Zou et al., 2024), illustrates how poisoned texts can lead to the generation of misinformation, revealing critical vulnerabilities in applications like cybersecurity and healthcare. However, these attacks often lack practicality as they concentrate on isolated factual queries. Such queries can often be countered through mitigation strategies such as reranking (De Stefano et al., 2024), which tend to overlook thematic or topic-level manipulation. Additionally, existing studies largely overlook controversial topics, where limited user understanding may increase susceptibility to opinion manipulation.
This paper addresses this critical gap by concentrating on the nuanced manipulation of opinions through adversarial attacks targeted at topics within RAG models, which presents a novel and urgent challenge given the increasing complexity and reliance on these systems.
In line with settings from previous data poisoning findings(Zou et al., 2024; Carlini et al., 2024), we consider a scenario where the attacker could inject a few carefully crafted poisoned texts into the knowledge base. For example, if the knowledge base includes texts sourced from Wikipedia, the adversary might inject poisoned content by maliciously editing Wikipedia entries. Besides, we focus on more practical and challenging black-box scenario, where no model information is disclosed, except that the attackers can query the target RAG and obtain responses, which contains the corresponding candidate referential documents (Chen et al., 2025).
In this paper, we propose the Topic-FlipRAG, which is a two-stage multi-granularity attack method designed to manipulating the stance polarity of RAGs with camouflaged modification of target documents. Specifically, in the first stage, we perform a stealthy adversarial modifications on target documents. This is achieved by incorporating the extensive internal general semantic knowledge in LLMs with its analysis and reasoning ability. In the second stage, inspired by adversarial semantic collision (Song et al., 2020), we utilize gradients from an open-sourced neural ranking model (NRM) to generate topic-specific adversarial triggers.
Our experiments show Topic-FlipRAG achieves 0.5 average stance variation (ASV) across four domains, significantly outperforming other baselines. It demonstrates such adversarial attacks can effectively manipulate RAG models to produce outputs aligned with specific topical stances, thus influencing how information is presented and perceived. Based on that, we also conduct the user experiments, which reveal a significant impact on users’ polarities toward controversial topics, with polarity shifts exceeding 16% after interacting with poisoned RAG systems. Futhermore, we explore several potential mitigation strategy, including perplexity-based detection, random masking, and reranking. Our results show these mitigations are inadequate against adversarial opinion manipulation attack, especially for Topic-FlipRAG, underscoring the necessity for novel mitigation strategies.
Our major contributions are as follows:
(1) We explore a novel and practical security scenario, Topic-oriented RAG Opinion Manipulation, which presents a broader and deeper threat to real-world users by enabling opinion manipulation across multiple topic-related queries.
(2) We propose Topic-FlipRAG, a knowledge-guided and multi-granularity black-box attack method tailored for topic-level opinion manipulation in RAG systems.
(3) Through extensive experiments, we demonstrate that Topic-FlipRAG achieves significantly better attack performance across all metrics compared to baseline methods. Furthermore, we evaluate its practical impact through user experiments, highlighting its potential to manipulate user opinions.
(4) We analyze the efficacy of several defense mechanisms and demonstrate their insufficiency in countering Topic-FlipRAG. This underscores the urgent need for robust and adaptive defense strategies against such attacks.
2. Background and Related Work
2.1. Retrieval Augmented Generation (RAG)
RAG models enhance their responses by accessing and incorporating external knowledge from large-scale databases or corpora during the generation process (Lewis et al., 2020; Shi et al., 2025; Liu et al., 2023a). By leveraging external data sources, RAG can provide more accurate and comprehensive answers, especially for queries requiring up-to-date information or specialized knowledge that may not be well-represented in the model’s training data (Siyue et al., 2024). Moreover, it can scale more effectively and flexibly by updating the retrieval corpus without necessitating extensive retraining of the generative component (Gao et al., 2023; Wu et al., 2024).
The workflow of a RAG systems is generally divided into two sequential phases: retrieval and generation. In the retrieval phase, upon the submission of a user query $q$, the RAG system retrieves $k$ relevant documents from the corpus $D$ with the highest embedding similarities to the query $q$. Specifically, for each document $d\in D$, the relevance score with the query $q$ is computed as $R(q,d)$. In this paper, we adopt the a more practical and widely used framework in LangChain that the retrieval phase is incorporated with historical-aware query rewriting and intention reasoning. In the generation phase, Given a the rewritten query $q^{\prime}$ , a set of top-k retrieved documents $D_{k}$, and access to the LLM, one can query the LLM with $q^{\prime}$ with the top-k retrieved documents $D_{k}$, the LLM generates an answer for original query $q$ by leveraging $D_{k}$ as context. This process is detailed in Appendix A.1.
2.2. Attacks to Retriever
Retriever vulnerabilities in RAG systems pose significant security risks due to their neural ranking dependencies (Goren et al., 2018). Adversarial attacks manipulate document rankings through semantic perturbations to promote target documents in query-specific results (Liu et al., 2022; Wu et al., 2023), undermining RAG’s core assumption of reliable high-ranked contexts. Current attacks are characterized along two axes: (1) Knowledge accessibility: White-box (full model access) vs. black-box (query-only) approaches. (2) Perturbation granularity: Word-level (Raval and Verma, 2020; Liu et al., 2023b, c; Wu et al., 2023), Phrase-level (Song et al., 2020; Liu et al., 2022; Chen et al., 2025), Sentence-level (Chen et al., 2023), and Hybrid (Liu et al., 2024).These neural semantic attacks parallel black-hat SEO tactics (Gyongyi and Garcia-Molina, 2005) but introduce novel challenges. Successful attacks propagate adversarial content to generators, enabling misinformation injection while evading traditional safeguards (Zhong et al., 2023; Chen et al., 2025), necessitating integrated security frameworks for RAG systems.
2.3. Attacks to LLMs
Existing attacks on LLMs include jailbreak attacks (Wei et al., 2024; Deng et al., 2024; Li et al., 2023; Lin et al., 2024), backdoor attacks (Kandpal et al., 2023; Lu et al., 2024; Cao et al., 2023), prompt injection (Liu et al., 2023; Perez and Ribeiro, 2022; Greshake et al., 2023), and poisoning attacks (Shafran et al., 2024; Zou et al., 2024; Cho et al., 2024; Zhong et al., 2023; Chen et al., 2025; Zhang et al., 2024). Poisoning attacks uniquely threaten RAG systems by injecting adversarial content into retrieval corpora to manipulate outputs. Current RAG poisoning studies focus on closed-domain factoid QA pairs (e.g., ”CEO of OpenAI”) (Zou et al., 2024; Zhong et al., 2023; Cho et al., 2024; Shafran et al., 2024; Zhang et al., 2024), overlooking the multi-query nature of real-world topic exploration(Liu et al., 2023c; Xue et al., 2024), e.g., ”smartwatch battery life” and ”health tracking accuracy” under the ”wearable tech” theme. This motivates our focus on practical topic-level universal perturbations mirroring universal adversarial examples.
Furthermore, existing defenses based on fact-checking are insufficient for opinion-based queries, such as ”Should genetic testing be regulated?”, which require nuanced reasoning rather than simple factual recall. Building on the insights from FlipedRAG (Chen et al., 2025), we tackle this significant gap by examining the adversarial manipulation of controversial topics. In such cases, LLMs are required to reason and synthesize multiple perspectives, making them particularly vulnerable to systematic knowledge poisoning.
3. Threat Model
Given a set of topic-queries $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{|\mathbf{Q}|}\}$, a RAG corpus $D$, and a desired target polarity $S_{t}$ (e.g., Pro or Con), The adversary’s objective is to steer the RAG system’s overall stance across the entire set of queries toward stance $S_{t}$. Let $\text{doc}_{\text{tar}}$ be the target document the attacker aims to promote. The black-box retrieval model RM assigns a relevance score $R(q,d)$ to each document $d\in D$ and returns the top-$k$ documents as $\text{RM}_{k}(q)=\{d_{1},d_{2},\ldots,d_{k}\}$. The LLM then generates an answer $(q,\text{RM}_{k}(q))$, and the extracted opinion from this response is $S_{o}=\text{S}(\text{LLM}(q,RM_{k}(q)))$.
Figure 2. An overview of our proposed Topic-Orientated Adversarial Opinion Manipulation Attack method for RAG systems.
3.1. Objective of the Adversary
The adversary seeks to modify the target document $doc_{tar}$ through a series of subtle perturbations $p_{adv}$, while preserving its original semantics and fluency, thereby transforming the corpus $D$ into $D^{\prime}=D(\text{doc};\text{doc}_{\text{tar}}\oplus p_{\text{adv}})$. By increasing the retrieval model’s relevance score $R(q,\text{doc}_{\text{tar}}\oplus p_{\text{adv}})$ for each query $q\in\mathbf{Q}$, the modified document becomes more likely to appear among the top-$k$ results $\text{RM}_{k}(q)$.
As a result, when the LLM generates answers conditioned on $(q,\text{RM}_{k}(q))$, the adversarially augmented document will influence the opinion $S(\text{LLM}(q,\text{RM}_{k}(q)))$ towards the target polarity $S_{t}$. Crucially, the objective extends beyond a single query: the adversary aims to ensure that, on average, the opinions produced by the LLM across the entire query set $\mathbf{Q}$ align with $S_{t}$. In other words, the attacker seeks:
$\max_{p_{\text{adv}}}\frac{1}{|\mathbf{Q}|}\sum_{q\in\mathbf{Q}}I\bigl{(}S(\text{LLM}(q,\text{RM}_{k}(q)))=S_{t}\bigr{)}\hskip 14.22636pt$
where $I(\cdot)$ is an indicator function that returns $1$ if the generated stance matches $S_{t}$ and 00 otherwise. By optimizing for this criterion, the adversary ensures that the aggregated stance across all topic queries shifts significantly toward the desired polarity.
3.2. Capabilities of the Adversary
In the black-box setting, the adversary is restricted to modifying a limited number of documents within the corpus $D$. Obtaining detailed information about the specific retrieval model employed by the target RAG system is typically impractical in real-world scenarios. The adversary has no access to the internal parameters or architecture of either the retrieval model or the large language model (LLM), and cannot alter the LLM’s prompt templates.
4. Methodology
4.1. Overview
In this study, we introduce Topic-FlipRAG, a two-stage, multi-granularity, topic-oriented attack pipeline designed to manipulate the stance polarity of retrieval-augmented generation (RAG) systems without altering the original content of the target document (Figure 2). Our proposed topic-oriented adversarial opinion manipulation attack for black-box RAG systems is also outlined in pseudo Algorithm 1. The pipeline proceeds as follows:
Input: Query list $Q$, target document $doc_{\text{tar}}$, target opinion $S_{t}$, iteration count $I$, keynode size $K$, augmentation factor $t$, sample size $N$, trigger length $T$, beam size $B$, top-$k$ size $K_{b}$, learning rate $\eta$
Output: Adversarial document $doc_{\text{adv}}$ (trigger + $doc_{\text{know}}$)
Phase 1: Knowledge-Guide-Attack
       INIT: $doc_{\text{selected}}\leftarrow\{\}$
       for $n=1\dots N$ do
             for $i=1\dots I$ do
                   keynodes $\leftarrow\mathrm{ExtractKeynodes}(Q,K)$
                   analysis $\leftarrow\mathrm{AnalyzeKeynodes}(keynodes,doc_{\text{tar}})$
                   adjustment $\leftarrow\mathrm{Adjust}(S_{t},\text{analysis},doc_{\text{tar}},t)$
[…]
                   filter_list $\leftarrow\mathrm{FilterDocument}(doc_{m},doc_{\text{tar}})$
[…]
            $t^{\prime}\leftarrow\mathrm{RewardingFunction}(t,\text{filter\_list},doc_{\text{tar}},Q)$
             $t\leftarrow t^{\prime}$
[…]
Phase 2: Adversarial Gradient-Based Trigger Generation
       INIT: $\mathbf{Z}\leftarrow[\mathbf{z}_{1},\ldots,\mathbf{z}_{T}],\quad,\text{where }\mathbf{Z}_{t}\in\mathbb{R}^{|\mathcal{V}|}\sim\mathcal{N}(0,1)$,
       $\quad\mathbf{z}_{t}\leftarrow w+\epsilon\;\text{for}\;w\in\mathbf{z}_{t}\;\text{if}\;w\in\mathcal{V}_{\text{sim}}$
[…]
      while not converged and iteration ¡ $I$ do
            $combined\_vec\leftarrow\mathrm{concat}(Z,doc_{\text{know}})$
             $\nabla_{Z}\leftarrow\mathrm{ComputeGradient}(combined\_vec,Q)$
             Update $Z$:
             $Z\leftarrow Z-\eta\cdot\nabla_{Z}$
             Perform Beam Search for the best trigger $Z$:
      $Trigger\leftarrow\mathrm{BeamSearch}(Z,doc_{\text{know}},T,B,K_{b})$
      OUTPUT: $doc_{\text{adv}}\leftarrow\mathrm{concat}(Trigger,doc_{\text{know}})$
       return $doc_{\text{adv}}$
Algorithm 1 Topic-FlipRAG(black-box)
Stage One: Knowledge-Guided Attack. In the initial phase, we leverage the extensive internal knowledge and reasoning capabilities of large language models to perform adversarial edits on the target document at a general semantic level. We identify and extract essential topic-related information nodes from the queries, then guide the LLM to integrate these nodes into $doc_{\text{tar}}$ through a multi-granular editing strategy across three semantic dimensions. Simultaneously, rigorous polarity control is applied to preserve the document’s original polarity, ensuring precise stance manipulation. A dynamic reward function refines each modification step, embedding critical elements and relevant information while maintaining minimal disruption to the document’s original semantics. This process establishes a firm basis for subsequent adversarial opinion manipulation.
Stage Two: Adversarial Trigger Generation. In the second phase, we enhance the similarity of the adversarially edited document $doc_{\text{know}}$ from Stage One. Leveraging an open-sourced Neural Ranking Model (NRM), we generate a concise adversarial trigger specifically crafted to maximize the alignment between $doc_{\text{know}}$ and the topic queries. This trigger is then fused with $doc_{\text{know}}$ and injected as a poisoned document into the RAG system’s database, significantly increasing its likelihood of appearing in retrieval results. Consequently, the system’s generated output shifts toward the desired stance polarity, thereby fulfilling the goal of topic-oriented adversarial manipulation.
4.2. Knowledge-Guided Attack
In this section, we introduce a three-phase framework, termed the knowledge-guided attack (abbreviated as know-attack), designed to systematically modify a target document $doc_{\text{tar}}$ to enhance its relevance to a given query set. The framework aims to achieve minimal modification of the document while effectively integrating topic-specific nodes, thereby influencing the final outputs of RAG systems. By operating across multiple levels of textual modification, this approach ensures semantic coherence, adheres to minimal editing principles, and maintains a predefined stance, facilitating precise and controlled manipulation of the target document.
Problem Statement. Given a target document $\text{doc}_{\text{tar}}$ and a query set $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{|\mathbf{Q}|}\}$, we seek a modified document $\text{doc}_{\text{know}}=\text{doc}_{\text{tar}}\oplus\mathcal{P}$ that increases its overall relevance to all queries while adhering to a target stance $S_{t}$.
$\max_{\mathcal{P}}\frac{1}{|\mathbf{Q}|}\sum_{q\in\mathbf{Q}}Sim(q,\text{doc}_{\text{know}})$
subject to:
$\|\mathcal{P}\|\leq\epsilon,\quad Sim(\text{doc}_{\text{know}},\text{doc}_{\text{tar}})\geq\lambda,\quad S(\text{doc}_{\text{know}})=S_{t}.$
Here, $\|\mathcal{P}\|$ constrains the degree of textual alteration to ensure minimal editing, and $Sim(\cdot)$ ensures sufficient semantic similarity between the original and modified documents. By incorporating these constraints, the resulting $\text{doc}_{\text{know}}$ not only aligns with the chosen stance $S_{t}$ and achieves higher relevance for all queries, but also maintains the document’s intrinsic coherence and readability, increasing its likelihood of influencing the RAG system’s outputs.
4.2.1. Phase 1: Key Node Extraction and Selection
Given a topic and its corresponding list of queries $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{|\mathbf{Q}|}\}$, we first leverage a large language model (e.g.GPT-4o-mini) to identify a set of $K$ key information nodes that are most salient for maximizing relevance with respect to the entire query list. These $K$ nodes form a candidate node list $\mathcal{L}$ encapsulating crucial topical aspects.
Next, we analyze $\text{doc}_{\text{tar}}$ to determine which nodes from $\mathcal{L}$ are underemphasized or entirely absent. The LLM generates a doc-specific node list with providing explicit reasoning for each missing or underemphasized node. This doc-specific list guides subsequent adversarial editing, ensuring that modifications focus on incorporating these critical nodes without introducing extraneous or tangential content.
4.2.2. Phase 2: Multi-Granular Adversarial Editing and Polarity Control
Inspired by (Liu et al., 2024) and adversarial strategies against black-box neural ranking systems (Liu et al., 2022; Wu et al., 2023), we employ a multi-granular editing approach leveraging the advanced language understanding of LLMs. The editing operates across three levels: (1) lexical substitutions, where individual words are replaced with synonyms or node-related terms while preserving original meaning; (2) sentential rewrites, which adjust sentence structures and introduce minimal relevant information to incorporate node details without altering semantics; and (3) phrase insertions, adding brief, contextually appropriate sentences containing node information at strategic points to maintain narrative coherence and content integrity.
During this adversarial editing process, we employ a module called Polarity Control to enforce a target stance $S_{t}$ constraint, guiding the document’s stance. For instance, if the target polarity is set to CON (opposing a viewpoint), the modifications introduced by the model consistently adopt negative or critical perspectives. This polarity control ensures that the integrated nodes and textual changes align with the desired stance, ultimately steering the RAG output towards the intended viewpoint.
4.2.3. Iteration with Rewarding Function and Final Output Selection
To adhere to the principle of minimal editing while ensuring effective document modification, we introduce an iterative framework guided by an augmentation factor $t$. This factor dynamically controls the extent and frequency of adversarial modifications during each editing iteration. A larger $t$ promotes more extensive and aggressive edits, while a smaller $t$ leads to a more conservative editing strategy. To mitigate fluctuations caused by the stochasticity of LLMs, we perform $I$ repeated sampling runs for each value of $t$. This ensures consistent performance by averaging the outcomes across multiple candidate modifications. After each editing iteration, $t$ is adaptively adjusted based on feedback from a filtering mechanism to maintain a balance between relevance and minimal editing.
In each iteration $n$ ($n=1,\dots,N$), for a given $t_{n}$, we generate $I$ candidate modified documents $\{\text{doc}_{m}^{(n,i)}\}_{i=1}^{I}$. Each candidate document is evaluated against two critical metrics:
Edit Distance(Edit Ratio): $d_{\mathrm{edit}}\bigl{(}\text{doc}_{m},\text{doc}_{tar}\bigr{)}$ measures the extent of textual modification compared to the original document.
Semantic Similarity: $d_{\mathrm{sem}}\bigl{(}\text{doc}_{m},\text{doc}_{tar}\bigr{)}$ evaluates how well the modified document preserves the original semantics.
To ensure minimal disruption, we specify a strict edit ratio threshold $\epsilon$ and a semantic similarity threshold $\theta$. A filtering indicator determines the validity of each candidate document:
$\mathrm{Filter}\bigl{(}\text{doc}_{m}\bigr{)}\;=\;\begin{cases}1,&\begin{matrix}\text{if }(d_{\mathrm{edit}}\bigl{(}\text{doc}_{m},\text{doc}_{tar}\bigr{)}\,\leq\,\epsilon)\\ \;\;\wedge\;(d_{\mathrm{sem}}\bigl{(}\text{doc}_{m},\text{doc}_{tar}\bigr{)}\,\leq\,\theta),\end{matrix}\\ 0,&\text{otherwise}.\end{cases}$
The augmentation factor $t$ is then dynamically updated based on the feedback from the filtering layer. To balance the trade-off between under-editing and over-editing, we introduce a partial edit threshold $\rho=0.75\,\epsilon$ and employ the following update rule:
$t_{n+1}\;=\;\begin{cases}t_{n}-\delta,&\text{if }\max\limits_{i}\mathrm{Filter}\bigl{(}\text{doc}_{m}^{(n,i)}\bigr{)}=0,\\[6.0pt] t_{n}+\delta,&\text{if }\exists\,\begin{matrix}(i\text{ s.t. }\mathrm{Filter}\bigl{(}\text{doc}_{m}^{(n,i)}\bigr{)}=1)\\ \;\wedge\;(d_{\mathrm{edit}}\!\bigl{(}\text{doc}_{m}^{(n,i)},\text{doc}_{tar}\bigr{)}<\rho),\end{matrix}\\ t_{n},&\text{otherwise}.\end{cases}$
Here, $\delta>0$ controls the aggressiveness of the adjustments. If no candidate satisfies the filtering constraints, $t$ is reduced to prevent excessive modifications. Conversely, if valid candidates exist but fail to meet the partial edit threshold $\rho$, $t$ is increased to encourage bolder edits. Otherwise, $t$ remains unchanged for stability.
After $N$ iterations, this process yields a set of $M$ valid candidate documents that satisfy the edit ratio and semantic similarity constraints. From this set, we apply a final selection phase to identify the optimal output. For each retained candidate $\text{doc}_{m}$, we use a neural ranking model (NRM) to estimate its relevance for each query $q_{i}\in\mathbf{Q}$, denoted as $R\bigl{(}q_{i},\text{doc}_{m}\bigr{)}$. The average relevance score for each candidate is computed as:
$\bar{R}\bigl{(}\text{doc}_{m}\bigr{)}\;=\;\frac{1}{|\mathbf{Q}|}\,\sum_{i=1}^{|\mathbf{Q}|}R\bigl{(}q_{i},\text{doc}_{m}\bigr{)}.$
The final modified document $\text{doc}_{\text{know}}$ is selected by maximizing the average relevance:
$\text{doc}_{\text{know}}\;=\;\arg\max_{\text{doc}_{m}}\;\bar{R}\bigl{(}\text{doc}_{m}\bigr{)}.$
The iterative process, directed by the rewarding function and filtering mechanisms, ensures that the final output $\text{doc}_{\text{know}}$ incorporates topic-specific nodes with minimal yet effective modifications. By maintaining semantic coherence and adhering to predefined constraints, it enhances the relevance of $\text{doc}_{\text{know}}$ to the query set while aligning its stance with $S_{t}$.
4.3. Adversarial Gradient-Based Trigger Generation
Building on the know-attack introduced in Section 4.2, which integrates general semantic-level node information into the target document $doc_{tar}$, we aim to further generate adversarial triggers targeting the retrieval model in the RAG system. Specifically, we append a short adversarial trigger $T$ to the previously constructed $\text{doc}_{\text{know}}$, resulting in the final adversarially augmented document $\text{doc}_{\text{adv}}$. Inspired by adversarial semantic collision (Song et al., 2020), we derive $T$ through a gradient-based optimization process using a public neural ranking model [^1] https://huggingface.co/nboost/pt-bert-base-uncased-msmarco. By integrating $\text{doc}_{\text{know}}$ with $T$, we align the document more closely with the query set $\mathbf{Q}$, increasing its prominence in the retrieval results. Once surfaced among the top-K documents, the strategically designed content within $\text{doc}_{\text{adv}}$ can direct the LLM’s final output toward $S_{t}$.
4.3.1. Problem Formulation
Our broader goal, as outlined previously in Section 3.2, is to influence the overall stance that the RAG system adopts when answering all queries in $\mathbf{Q}$. While our ultimate objective is to ensure that the LLM’s output consistently aligns with the target polarity $S_{t}$ across the entire query set, directly optimizing this stance-level objective remains challenging due to the lack of gradient signals from the LLM.
To address this challenge, we first concentrate on an intermediate relevance-based objective. Let $RM(q_{i},\text{doc})$ represent the neural ranking model’s relevance score for a query-document pair $(q_{i},\text{doc})$. Given a modified document $\text{doc}_{\text{adv}}=[\text{doc}_{\text{know}},T]$, our interim goal is to maximize the average relevance score in all queries:
$\max_{T}\frac{1}{|\mathbf{Q}|}\sum_{i=1}^{|\mathbf{Q}|}RM\bigl{(}q_{i},[\text{doc}_{\text{know}},T]\bigr{)}.$
By enhancing the document’s relevance to each $q\in\mathbf{Q}$, we increase the likelihood that $\text{doc}_{\text{adv}}$ appears among the top-$k$ retrieved documents for all queries. Achieving this intermediate objective paves the way for influencing the stance of the RAG system’s responses at a broader, topic-queries scale.
4.3.2. Gradient-Based Optimization
We begin by leveraging gradients from a neural ranking model over the entire topic-query set, initially refining a soft trigger representation. Next, we employ beam search to transform the refined soft trigger into an optimal discrete token sequence. By iterating between gradient-based adjustments and beam search selection, the approach converges on a trigger that maximizes the document’s relevance across all topic queries. The following steps detail the core components of this optimization procedure.
Gradient Optimization: During each optimization step, we compute and combine gradients for all queries in $\mathbf{Q}$. Additionally, a word enhancement argument is introduced, which incorporates query-specific keywords and lexical distributions from the model’s vocabulary. This guided selection of candidate tokens captures semantic overlaps among multiple queries and enhances overall trigger relevance.
Beam Search: In the beam search phase, we evaluate candidate triggers—along with doc-specific augmentation—to select those that maximize combined similarity scores across the entire query set. This ensures the chosen trigger reliably improves the target passage’s ranking impact when integrated with $\text{doc}_{\text{know}}$.
Doc-Specific Augmentation: At each iteration of gradient calculation and beam search, the trigger’s intermediate soft representation is evaluated only when concatenated with $\text{doc}_{\text{know}}$ and considered jointly with all queries. Although this increases computational overhead, it ensures tight alignment with the target document’s context, yielding more pronounced improvements in ranking performance.
4.3.3. Final Output: $\text{doc}_{\text{adv}}$ and RAG Opinion Influence
After convergence, the adversarial trigger $T$ is appended to $\text{doc}_{\text{know}}$ to form the final adversarial document:
$\text{doc}_{\text{adv}}=[\text{doc}_{\text{know}},T].$
This $\text{doc}_{\text{adv}}$ is then injected into the RAG system’s corpus $D^{\prime}$, significantly increasing its likelihood of being retrieved among the top-$k$ documents for each query $q\in Q$. Consequently, $\text{doc}_{\text{adv}}$ frequently appears in retrieval results, leading the LLM to generate outputs that align with the desired stance $S_{t}$. This process effectively manipulates the overall stance of the RAG system by ensuring that the augmented document is both prominently retrieved and influences the generated opinions towards $S_{t}$.
5. Experiments
5.1. Datasets
MSMARCO. We utilized the MS MARCO Passage Ranking dataset as the data source to evaluate the ability of our method to improve document rankings under more challenging topic-query tasks. Specifically, we assessed whether our method could significantly enhance the ranking of documents by the retrieval model within a RAG system.
To construct topic-lists for evaluation, we applied a K-means clustering algorithm to group similar queries, forming topics that each contained a series of related queries. To further evaluate the performance of our method under extreme topic-query scenarios, we applied an intra-topic similarity filtering process. Only topics with queries exhibiting high semantic diversity and containing a sufficient number of queries were retained.
This process resulted in 29 topics, with each topic containing an average of 22.28 queries. The average similarity score within each topic was approximately 0.5, indicating sufficient diversity among queries to ensure a rigorous evaluation. This curated dataset enabled us to test the robustness of our method in handling highly diverse and challenging topic-query tasks within a RAG system.
PROCON. To conduct our experiments, we utilized controversial topic data scraped from the PROCON.ORG website. This dataset includes over 80 topics covering various domains, such as society, health, government, and education. Each topic is discussed from two stance labels {PRO (support), CON (oppose)}, with passages arguing from these perspectives.
To simulate real-world user interactions with a RAG system, we instructed a large language model (GPT-4o) to act as a user and generate 40 potential sub-queries for each topic. These sub-queries were designed to reflect the diverse questions and concerns users might raise when exploring a specific controversial topic.
After generating the sub-queries, we applied a similarity filtering process to ensure diversity by retaining only those with a similarity score below approximately 0.85. The filtering step effectively removed redundant queries while preserving a wide range of perspectives. As a result, the final set of topic-queries achieved an average similarity score of approximately 0.7, ensuring that the queries were sufficiently diverse yet semantically relevant. This process provided a robust and balanced sub-queries set for evaluation.
5.2. Experiment Details
The specific setting details for the Topic-queries RAG manipulation experiment are as follows:
(1) Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. The LLMs adopted in RAG are the open-source models Meta-Llama-3.1-8B-Instruct (Llama3.1), Qwen-2.5-7B-Instruct(Qwen2.5). The system prompt and additional detailed descriptions are provided in Appendix A.1.
(2) Retrieval Model Specification. We benchmark three dominant dense retrievers—Contriever (Gao and Callan, 2021), DPR (Karpukhin et al., 2020), and ANCE (Xiong et al., 2020).By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking.
(3) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.
(4) Experimental parameters. In knowledge-guided attack process, we set the maximum editing distance $\epsilon$ to 0.2, the semantic similarity threshold $\lambda$ to 0.85, and the number of iterations $N$ to 5. For adversarial trigger generation, we used a beam size of 3, top-$k$ of 10, a batch size of 32, a temperature of 1.0, a learning rate of 0.005, and a sequence length of 10. In RAGblack, $k$ (the number of retrieved documents) is set to 3, with the LLM temperature also fixed at 1.0 to mirror real-world conditions.
(5) Poisoned Target. For the PROCON dataset, to investigate the manipulation performance under more challenging conditions, we performed relevance ranking for the documents with respect to each topic-query set $Q$ and the target stance $S_{t}$ From the ranked list, we selected the last five documents (i.e., those with the lowest relevance) as the target poisoned documents. For the MS MARCO dataset, we utilized the top-1000 relevance-ranked passage list for each topic-query set. From this list, we selected the passage with the lowest average rank as the target passage. This approach ensures that the evaluation focuses on passages that are least relevant to the target queries, thus providing a more rigorous benchmark for the proposed method.
(6) Experimental environment. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
5.3. Research Questions
We propose four research questions to evaluate the effectiveness of our method in the topic-queries task, focusing on black-box NRM attacks and opinion manipulation to RAGs.
RQ1: Can Topic-FlipRAG significantly enhance the rankings of target documents in the RAG retriever for topic-queries?
RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
RQ3: Does topic-oriented opinion manipulation significantly impact users’ perceptions of controversial topics?
RQ4: How robust does Topic-FlipRAG against exisiting mitigation mechanism?
5.4. Baseline Settings
To assess the effectiveness of our proposed method, we compare it against adversarial attack baselines designed for black-box, topic-oriented RAG scenarios, ensuring minimal modifications to the original documents. We exclude BadRAG(Xue et al., 2024), a backdoor RAG attack limited to white-box scenarios, and topic-IR-attack(Liu et al., 2023c), as its incomplete implementation prevents reliable replication. For the selected baseline methods, we adapted them to meet the requirements of our task while preserving their core components. A brief overview of the baseline methods is provided below, with detailed descriptions available in Appendix A.2.
PoisonedRAG. Zou et al.(Zou et al., 2024) propose an approach adaptable to both black-box and white-box settings. For our task, we employ its black-box strategy by inserting a randomly chosen query from the topic-queries set $Q$ at the beginning of each document.
PAT. This gradient-based adversarial retrieval attack uses a pairwise loss function to generate triggers that meet fluency and coherence constraints. We adapt PAT to produce triggers $T_{\text{pat}}$ for target documents within the topic-queries set, evaluating their effectiveness under black-box conditions.
Collision. This method generates adversarial paragraphs (collisions) via gradient-based optimization to produce content semantically aligned with the target query. In a topic-queries context, we create collisions for the entire topic-queries set and examine their transfer performance on black-box RAG retrievers.
These baseline methods provide benchmarks for comparing the efficacy of our approach in a fully black-box, topic-oriented RAG attack scenario.
5.5. Evaluation Metrics
[…]
Ranking Attack Success Rate (RASR). RASR captures how frequently target documents are successfully boosted in each query’s ranking. Higher values indicate greater attack effectiveness.
Boost Rank (BRank). BRank is the average rank improvement for all target documents under each query. A target document contributes negatively if its rank is unintentionally lowered.
Top-50, Top-500. These metrics represent the percentage of target documents that move into specific ranking thresholds in the MS MARCO Dataset after manipulation. Higher percentages imply more effective promotion of target documents.
For RQ2, we employ Average Stance Variation (ASV) to assess how significantly our opinion manipulation influences the LLM’s responses in a black-box RAG. To address the natural variability of controversial topics and the inherent instability of large language models, we also propose Real Adjusted ASV ($\Delta$-ASV).
Average Stance Variation (ASV). ASV is defined as the absolute difference between the manipulated opinion score and the original opinion score assigned to an LLM response (0 = opposing, 1 = neutral, 2 = supporting). A higher ASV signifies a more pronounced shift in polarity and hence greater manipulation effectiveness.
[…]
6. Experimental Results Analysis
6.1. RQ1: Can our method significantly enhance the rankings of target documents in the RAG retriever for topic-queries?
To explore RQ1, we first conducted white-box IR manipulation experiments on MS MARCO and PROCON datasets. The results of these experiments are presented in Table 1 and Table 2, respectively.
MS MARCO Dataset Validation. The results demonstrate that Topic-FlipRAG significantly outperforms all other baselines across all evaluation metrics, even under challenging conditions where the MS MARCO dataset was configured with reduced internal similarity among topic-queries to simulate more extreme scenarios. On this dataset (Table 1), Topic-FlipRAG achieves an almost perfect Retrieval Attack Success Rate (RASR) of 99.29%, substantially higher than Poisoned-RAG (64.39%), Collision (93.19%), and PAT (89.63%). Furthermore, it achieves the highest B-rank (521.30) and superior coverage of top-ranked documents, with 18.35% for top-50% and 79.49% for top-500% metrics. These results highlight Topic-FlipRAG’s ability to effectively manipulate rankings across the entire topic-queries set, achieving significant ranking enhancements for each query while simultaneously targeting multiple queries with high success rates.
Table 1. White-box IR attack performance (RASR, Top-50, and Top-500 are reported in %) on MS MARCO dataset. Bold indicates the best attack performance.
Method RASR B-rank Top-50 Top-500
Collision 93.19 245.19 0.46 40.56
PAT 89.63 208.06 0.31 33.90
PoisonedRAG 64.39 73.52 2.94 11.15
Topic-FlipRAG 99.29 521.30 18.35 79.49
Table 2. White-box and Black-box IR attack results (%) on PROCON dataset. Bold indicates the best attack performance.
Setting White-box Black-box
Target IR model Target Stance BERT Contriever ANCE DPR
Method RASR top3-v RASR top3-v RASR top3-v RASR top3-v
Collision CON 48.44 23.23 21.27 8.38 18.79 10.41 16.87 5.97
PRO 43.64 20.85 21.44 8.24 15.97 7.73 14.49 5.23
PAT CON 37.85 17.23 16.40 7.15 19.51 12.36 13.99 5.65
PRO 31.99 13.35 13.07 4.30 13.44 9.14 8.61 2.26
PoisonedRAG CON 39.87 18.64 25.52 11.99 34.10 14.91 31.25 14.13
PRO 32.75 14.04 22.04 8.71 25.37 9.81 21.26 8.31
Topic-FlipRAG CON 79.42 44.71 51.90 25.57 45.69 20.12 42.27 18.83
PRO 75.01 37.04 49.20 21.55 42.12 20.57 41.84 16.41
PROCON Dataset Validation. On the PROCON dataset (Table 2), the results strongly confirm the superiority of Topic-FlipRAG. In the white-box setting, for both target stances (CON and PRO), Topic-FlipRAG achieves the highest RASR (0.79 for CON, 0.75 for PRO) and the largest top3-v (0.45 for CON, 0.37 for PRO). In contrast, Poisoned-RAG, which uses a simple query+ strategy, shows considerably lower effectiveness (RASR: 0.40 for CON, 0.33 for PRO), illustrating the limitations of its approach for topic-oriented attacks. While the gradient-based methods Collision and PAT match or exceed Poisoned-RAG in opinion manipulation, they still trail Topic-FlipRAG. This outcome highlights Topic-FlipRAG’s ability to conduct effective manipulation across entire sets of topic-queries.
To further evaluate our method’s practical effectiveness, we conducted black-box IR manipulation experiments on the PROCON dataset using three widely adopted dense retrievers (Contriever, ANCE, DPR). Owing to the inherent robustness of these retrievers, performance declined for all attack methods compared to the white-box setting. Nevertheless, Topic-FlipRAG displays the most consistent performance, retaining relatively high manipulation effectiveness across all models and metrics. For instance, under Contriever, Topic-FlipRAG achieves the highest RASR (0.52 for CON, 0.49 for PRO) and top3-v (0.26 for CON, 0.22 for PRO), notably outperforming Poisoned-RAG, Collision, and PAT. A similar pattern emerges for ANCE and DPR, where Topic-FlipRAG continues to hold a clear advantage.
Topic-FlipRAG’s superior performance in transfer attack settings can be attributed to its integration of the LLM’s general semantic knowledge and NRM gradient guidance, along with a design specifically tailored for the topic-queries context. This combination enables the method to generate adversarial triggers that are both generalizable and contextually precise, allowing it to adapt effectively to varying retriever architectures without relying on the target IR model’s parameters. In contrast, gradient-based methods like Collision and PAT depend heavily on gradient information from the target model, leading to substantial performance degradation in transfer attack scenarios. For instance, Collision achieves a maximum RASR of only 0.21 on Contriever for both CON and PRO stances, while PAT performs even worse, with RASR values of 0.16 (CON) and 0.13 (PRO).
Additionally, our findings reveal varying levels of vulnerability among dense retrievers. Contriever appears the most susceptible (exhibiting higher RASR and top3-v scores for all attackers), whereas DPR is notably more resilient (e.g., RASR for Topic-FlipRAG falls to 0.42). These observations underscore the importance of careful retriever selection when designing RAG systems to mitigate adversarial risks.
Table 3. Topic-Oriented RAG attack results on PROCON dataset, including ASV on different domains (Health & Environment, Government & Politics, Education, and Society & Culture), average ASV, and $\Delta$ASV. Bold shows the best performances.
Method Target Stance Contriever + qwen-2.5-8B-instruct Contriever + llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
Clean CON 0.21 0.27 0.24 0.23 0.24 0.22 0.26 0.18 0.27 0.24
PRO 0.21 0.26 0.24 0.24 0.24 0.23 0.27 0.21 0.27 0.25
Collision CON 0.30 0.36 0.35 0.34 0.34 0.10 0.30 0.35 0.23 0.27 0.30 0.06
PRO 0.27 0.40 0.30 0.28 0.32 0.08 0.29 0.37 0.24 0.28 0.31 0.06
PAT CON 0.30 0.27 0.28 0.27 0.28 0.04 0.26 0.35 0.25 0.32 0.30 0.06
PRO 0.29 0.25 0.31 0.33 0.29 0.05 0.28 0.30 0.30 0.30 0.29 0.04
PoisonedRAG CON 0.30 0.35 0.26 0.38 0.33 0.09 0.34 0.39 0.31 0.46 0.38 0.14
PRO 0.27 0.41 0.28 0.32 0.34 0.10 0.31 0.38 0.29 0.39 0.35 0.10
Topic-FlipRAG CON 0.51 0.46 0.48 0.55 0.49 0.25 0.54 0.51 0.30 0.64 0.50 0.26
PRO 0.43 0.55 0.42 0.51 0.49 0.25 0.44 0.58 0.41 0.44 0.48 0.24
6.2. RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
Figure 3. Case study of Topic-oriented Opinion Manipulation On Topic: ”Is Binge-Watching Good for You?”
Table 3 presents the results of different methods on RAG output opinion manipulation, comparing the performance of two retrieval-augmented generation (RAG) systems: contriever + qwen-2.5-8B-instruct and contriever + llama3.1-8B-instruct. Additionally, experiments were conducted using alternative retrievers such as ANCE and DPR, with results provided in Table8 and Table 9 (in Appendix).
To evaluate the manipulation effectiveness across different topic domains, the PROCON dataset was divided into four subsets: Health & Environment (H&E), Government & Politics (G&P), Education (EDU), and Society & Culture (S&C). Each reported avg-ASV reflects the average manipulated stance polarity (PRO or CON) of the RAG outputs across all domains, while avg-$\Delta$-ASV represents the manipulation effect, computed by subtracting the clean-state avg-ASV (clean) from the manipulated avg-ASV. Both ASV and $\Delta$-ASV range from 0 to 2, where higher values indicate stronger manipulation effects.
The results demonstrate that Topic-FlipRAG achieves significantly better manipulation across all domains compared to baseline methods, including PoisonedRAG. Specifically, Topic-FlipRAG achieves an avg-ASV of 0.49 for qwen-2.5 and 0.48 for llama3.1, outperforming all other methods. PoisonedRAG, the second most effective method, achieves avg-ASV values of 0.33 for qwen-2.5 and 0.35 for llama3.1, while other baselines such as Collision and PAT remain within the range of 0.28–0.31. In terms of manipulation effect, measured by avg-$\Delta$-ASV, Topic-FlipRAG also achieves the highest values: 0.26 for qwen-2.5 and 0.24 for llama3.1. These results highlight the robust manipulation capability of Topic-FlipRAG, surpassing both direct adversarial document injection methods like PoisonedRAG (0.09 for qwen-2.5 and 0.10 for llama3.1) and stance-optimization methods such as and PAT (0.06 for both LLMs).
Table 3 and the appendix results further reveal notable differences between LLMs and retrievers regarding susceptibility to topic-oriented opinion manipulation. Overall, llama3.1 exhibits higher avg-ASV and avg-$\Delta$-ASV across all baseline methods compared to qwen-2.5, indicating that llama3.1 is more vulnerable to adversarial documents promoted to the top-3 of the ranking. This suggests that llama3.1 tends to generate responses that are more aligned with adversarial stances in the context. Similarly, among retrievers, contriever-based RAG systems achieve higher avg-$\Delta$-ASV values compared to ANCE and DPR, as detailed in the appendix, aligning with RQ1 findings that contriever produces superior ranking distortion, particularly for top-3 results.
The manipulation effectiveness of Topic-FlipRAG also varies across different topic domains. For instance, for llama3.1, Society & Culture (S&C) questions are the most susceptible to opposing stance manipulation (avg-ASV = 0.55), while Education (EDU) questions are relatively resistant (avg-ASV = 0.41). Health & Environment (H&E) questions demonstrate greater difficulty in manipulating supporting stances (PRO) compared to opposing stances (CON). Both qwen-2.5 and llama3.1 display higher avg-ASV values for supporting stance manipulation in Government & Politics (G&P), while Education (EDU) emerges as the most challenging domain for stance manipulation across both LLMs.
In order to illustrate the manipulation results more vividly, we randomly selected the topic “Is Binge-Watching Good for You?” for a case study, with the findings presented in Figure 3. As demonstrated by the distribution variation of RAG response opinions in the figure,, Topic-FlipRAG can manipulate RAG-generated content across multiple queries within the specific topic, shifting the overall opinion distribution to match a specified stance. Initially, responses on this topic span a mix of supportive, opposing, and neutral perspectives; following manipulation, the distribution becomes predominantly supportive when the target stance is “Pro” (blue) or predominantly opposing when the target stance is “Con” (red).
Figure 3 further demonstrates Topic-FlipRAG’s impact on multiple sub-queries. For instance, in “Binge-watching: addictive or relaxing?”, the unmanipulated response presents a balanced view. However, once manipulated, it adopts a distinctly polarized stance—emphasizing the relaxing aspects under “Pro,” or highlighting detrimental, addictive effects under “Con.” Similarly, for “Opinions on binge-watching trends?”, the original responses encompass diverse perspectives, whereas the manipulated version exhibits a consistent bias aligned with the designated stance. These examples underscore Topic-FlipRAG’s capacity to impose coherent opinion shifts across multiple queries under a single topic.
6.3. RQ3: Does topic-oriented opinion manipulation significantly impact users’ perceptions of controversial topics?
Figure 4. Empirical comparative experiments of adversarial opinion manipulation attacks on user cognition.
To explore the effect of opinion manipulation on users’ perceptions of controversial topics, we conducted a user study involving 54 college students (29 males, 25 females) to assess how opinion manipulation influences perceptions of controversial topics. Participants were split into two equal groups (Group-clean and Group-poisoned, 27 members each), remaining unaware of the study’s true purpose and operating independently. A simple QA service based on RAG was implemented, and Topic-FlipRAG was employed to poison documents with specific stances on two designated controversial topics. For one topic, supporting documents were poisoned; for the other, opposing documents were poisoned. Each participant interacted with the QA system over three rounds per topic, mirroring real-world usage patterns, and subsequently rated their stance on a 7-point Likert scale (normalized to 0 for strong opposition, 1 for strong support).
Figure 4 illustrates the significant shifts in user opinions resulting from Topic-FlipRAG. For Topic 1 (Should People Become Vegetarian?), the manipulated stance was PRO. Group-clean exhibited a dispersed distribution with an average score of 0.45, reflecting diverse views on this controversial issue. By contrast, Group-poisoned recorded a notably higher mean score of 0.61, indicating successful manipulation toward pro-vegetarian sentiment. For Topic 2 (Should Humans Colonize Space?), the stance was CON. Although Group-clean originally showed strong support (average score 0.76), Group-poisoned achieved a reduction to 0.59, effectively nudging opinions toward neutrality.
These findings underscore the potency of Topic-FlipRAG in altering user perceptions via RAG, whether opinions are initially neutral or polarized. The consistent polarity shifts across both topics reveal the potential risks posed by adversarial attacks in shaping public discourse on contentious issues.
Table 4. Ablation study (top3-ori, top3-att, and top3-v are reported in %) on the PROCON dataset. w/o denotes “without”.
Method IR Llama3.1 Qwen2.5
top3-ori top3-att top3-v ASV $\Delta$ ASV ASV $\Delta$ ASV
Topic-FlipRAG 42.83 70.24 27.41 0.64 0.37 0.55 0.32
w/o adv-trigger 42.83 51.12 8.29 0.37 0.10 0.38 0.15
w/o know-attack 42.83 59.40 16.57 0.40 0.12 0.41 0.18
clean 0.27 0.23
6.4. Ablation Study
Effectiveness of Core Components of Topic-FlipRAG. We investigated the individual contributions of the two main components of Topic-FlipRAG: know-attack and adversarial trigger generation (referred to as adv-trigger). As shown in Table 4, both components play significant roles in achieving the overall attack effectiveness, albeit with varying impacts on different aspects of the task.
Adv-trigger and know-attack complement each other by leveraging distinct strengths to enhance the overall effectiveness of the Topic-FlipRAG framework. Adv-trigger, through its gradient-based generation process, directly optimizes the retriever mechanism, resulting in significant ranking improvements for adversarially targeted documents. In contrast, know-attack, while less impactful on ranking optimization, utilizes general semantic knowledge to achieve precise and consistent polarity manipulation in LLM outputs, reinforcing the adversarial influence on generated responses. This synergy allows for more comprehensive and precise control over the RAG framework’s behavior, thereby maximizing the overall efficacy of the attack strategy.
Impact of Rewarding Function in Topic-FlipRAG. The rewarding function in Topic-FlipRAG dynamically adjusts the augment factor $t$ in the know-attack process (as described in Algorithm 1), controlling the aggressiveness of edits in each iteration. By leveraging semantic similarity and edit distance filtering, the rewarding function ensures an adaptive improvement strategy, enhancing the overall performance.
As shown in Table 5, the rewarding function significantly impacts the effectiveness of know-attack in IR manipulation. Compared to fixed augment factor $t$ values, the dynamic adjustment achieves higher attack breadth (evidenced by increased RASR) and depth (higher top3-v scores). Table 6 further highlights its influence on RAG opinion manipulation, where the dynamic approach yields higher ASV and $\Delta$ ASV compared to fixed $t$ settings. Notably, the rewarding function’s influence on RAG output polarity manipulation surpasses its impact on IR manipulation. This can be attributed to its role in improving the quality of generated $\text{doc}_{\text{know}}$, where quality encompasses both enhanced ranking boosts (related to IR manipulation) and stronger polarity reinforcement. These findings underscore the rewarding function’s critical contribution to optimizing the know-attack strategy for both ranking and opinion manipulation objectives.
Table 5. Comparative analyses of rewarding function and fixed augment factor $t$ on IR manipulation attacks (%).
RASR top3-ori top3-att top3-v
know-attack 22.25 42.83 52.12 8.29
t=1 17.42 42.83 49.28 6.45
t=2 16.83 42.83 49.73 6.89
t=3 18.54 42.83 49.02 6.18
Table 6. Comparative analyses of rewarding function and fixed augment factor $t$ on RAG opinion manipulation.
Setting Topic-FlipRAG t=1 t=2 t=3
ASV 0.64 0.58 0.56 0.59
$\Delta$ ASV 0.37 0.31 0.29 0.32
Impact of Polarity Control in Topic-FlipRAG. Figure 5 illustrates the effect of the Polarity Control component on the sentiment of adversarial documents ($\text{doc}_{\text{adv}}$) when the target polarity is set to CON. Negative sentiment scores represent negative polarity, while positive scores indicate positive polarity. (For PRO, see Figure 10 in the Appendix.) The results show that, without Polarity Control, $\text{doc}_{\text{adv}}$ gravitates toward a neutral polarity. With Polarity Control enabled, Topic-FlipRAG effectively aligns $\text{doc}_{\text{adv}}$ with the desired target polarity (CON).To further investigate the influence of attacks on document sentiment, we present measurement results in Figure 12 and Figure 13 in the Appendix.
To further investigate how Polarity Control influences RAG opinion manipulation, we examine not only the sentiment but also the stance of $\text{doc}_{\text{adv}}$ (i.e., its degree of alignment with the target polarity). For both PRO and CON settings, we randomly selected five topics and measured stance (normalized from $-1$ to $1$, where $-1$ signifies strong opposition and $1$ signifies strong support). As shown in Figure 11 (in the Appendix), without Polarity Control, the stance of $\text{doc}_{\text{adv}}$ shifts closer to neutrality, weakening its polarity. In contrast, incorporating Polarity Control preserves stance, enabling more effective manipulation of the LLM’s final outputs.
Figure 5. Impact of Polarity Control module on documents sentiment. w/o denotes “without”. (target polarity:CON).
The notable deviation in $\text{doc}_{\text{adv}}$ polarity without Polarity Control arises from the nature of know-attack: while it leverages the internal knowledge and reasoning capabilities of LLMs, it also inherits inherent polarity biases. Absent strict polarity enforcement during each editing step, the generated $\text{doc}_{\text{adv}}$ tends to drift from the intended polarity. By enforcing the desired polarity at every phase of know-attack, the Polarity Control component supports a more consistent and effective manipulation of RAG polarity outcomes.
Figure 6. Log perplexity (PPL) distributions on the PROCON dataset.
Potential Effectiveness by Altering Poisoning Target. Table 7 illustrates the impact of different poisoned methods on IR and RAG manipulation results. In our prior experiments, we selected the least relevant five documents (last5) as the poisoning targets to evaluate Topic-FlipRAG’s performance in a challenging manipulation scenario. This design aimed to test the robustness of the attack under more constrained conditions. However, in real-world attack scenarios, an attacker might reasonably choose the topically most relevant documents (top5) as poisoned targets to achieve more effective opinion manipulation.
The experimental results show that poisoning the top5 documents yields improved manipulation performance, as evidenced by higher top3-v and ASV scores. Specifically, poisoning the top5 documents achieves a top3-v score of 0.37 and an ASV score of 0.71, both surpassing the performance observed with the last5 method. This indicates that Topic-FlipedRAG could cause even greater harm if deployed in real-world scenarios, where attackers have the flexibility to select highly relevant documents as targets. These findings underscore the pressing need to develop robust defenses against such advanced adversarial strategies.
Table 7. Performance comparison of poisoning target on the PROCON dataset.(top3-ori, top3-att, top3-v are reported in %.)
Poisoning Target IR Llama3.1
top3-ori top3-att top3-v ASV $\Delta$ ASV
last5 42.63 70.24 27.61 0.64 0.37
top5 42.63 79.18 36.55 0.72 0.44
clean 0.27
6.5. RQ4: How robust does Topic-FlipRAG against exisiting mitigation mechanism?
Mitigation by perplexity. Perplexity (PPL) is widely used to measure the quality of texts, which is also utilized to defend against attacks to LLMs and IR model(Jain et al., 2023; Gonen et al., 2022; Liu et al., 2022). A large perplexity of a text means it is of low quality. We utilize perplexity to detect malicious texts.
Figure 6 shows the log perplexity (PPL) distributions evaluated using GPT-2 (Radford et al., 2019) on PROCON dataset documents manipulated by various attack methods. The results reveal the limitations of perplexity-based defenses in distinguishing poisoned documents from clean ones, as significant overlaps in PPL distributions persist. This overlap makes it difficult to set effective thresholds: strict thresholds lead to high false positive rates, while loose thresholds fail to detect attacks, undermining the defense’s efficacy.
Specifically,Collision exhibits the largest deviation in PPL distribution, making it relatively easier to detect. In contrast, Topic-FlipRAG and PAT produce poisoned documents with distributions closely resembling the original dataset, posing significant challenges for detection. PoisonedRAG, which lacks gradient-guided optimization, shows almost no shift in PPL, further highlighting the insufficiency of perplexity-based filtering for nuanced manipulations.
Figure 7. Attack performance of different random mask rate.
Mitigation by Random Mask. Random Mask is a robustness enhancement method that randomly masks a proportion of tokens with a placeholder token $m$ during embedding (Zeng et al., 2023). By averaging predictions over multiple masked input versions, it reduces sensitivity to small perturbations and enhances model stability. In our implementation, we averaged outputs over three masked copies of each input to assess performance under varying mask rates.
Figure 7 presents experimental results on the PROCON-Dataset targeting the CON stance. When the mask rate is below 0.3, Topic-FlipRAG achieves notably higher manipulation metrics (RASR ¿ 55%, top3-v ¿ 20%) compared to other baselines, underscoring its robust manipulation capability. As the mask rate increases, the effectiveness of all methods declines; however, Topic-FlipRAG still maintains the highest attack success among the baselines. Nonetheless, excessively large mask rates also degrade normal RAG performance, revealing the trade-off between heightened defense and overall system functionality. Consequently, Random Mask alone is insufficient against sophisticated attacks like Topic-FlipRAG, indicating that additional measures are required to strengthen system robustness without undermining functionality.
Mitigation by reranking. We investigate a reranking mechanism within the RAG framework to defend against adversarial attacks. This approach uses a different IR model to refine and reorder initial retrieval results, thus diminishing attack efficacy. The defense rationale is twofold: (1) existing attacks typically rely on simulating the target IR model (Cho et al., 2024; Liu et al., 2022, 2024), and introducing a distinct reranker weakens this dependency; (2) varying IR models exhibit different robustness levels, and reranking adds resilience.
We tested this method on the PROCON-Dataset with the stance CON, using Contriever for retrieval and DPR for reranking. As illustrated in Figure 8, most adversarial methods show a slight decline in manipulation effectiveness under reranking, except for PoisonedRAG, which is less sensitive due to its simpler query-plus strategy. In contrast, Collision, PAT, and Topic-FlipRAG depend on gradient information from open-sourced IR models, leading to varied transferability when attacking a different black-box IR pipeline. Despite the overall decrease in attack performance, Topic-FlipRAG maintains a relatively high success rate, demonstrating that reranking alone offers only partial defense against sophisticated adversarial techniques.
Figure 8. Comparative analysis of the performance metrics (RASR and top3-v) across different baselines before and after reranking mitigation.
7. Conclusion
This study highlights the critical vulnerability of RAG systems to sophisticated opinion manipulation attacks in real-world scenarios. By shifting focus from traditional factual or single-query attacks to topic-oriented adversarial manipulations, we reveal how attackers can systematically exploit the knowledge synthesis capabilities of LLMs to alter information perception and propagate biased narratives. The proposed Topic-FlipRAG framework demonstrates that combining adversarial ranking perturbations with LLMs’ inherent reasoning abilities enables effective stance polarity shifts across interconnected queries, even in complex multi-perspective contexts. Experiments show that existing mitigation strategies inadequately address semantically sophisticated attacks. In the future, we will explore enhanced transparency protocols to maintain RAG systems as reliable information mediators.
Appendix A Appendix
A.1. Full experiment details
(1)Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. Based on a retriever and a large language model (LLM), Conversational RAG mainly consists of a history-awaring retriever to identify users’ intention and a question-answering chain to produce output, as shown in Figure9. The LLMs adopted are the open-source models Meta-Llama-3.1-8B-Instruct (Llama3.1), Qwen-2.5-7B-Instruct(Qwen2.5). They perform well across various tasks among all open-source LLMs. The prompt connecting the retriever and the LLM in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain:
Use the following pieces of retrieved context to answer the question. Keep the answer concise. Context: {context}. Question: {question}.
Figure 9. The framework of Conversational RAG.
(2) Target retriever model in RAG-system. People usually use dense retrieval models as the retriever in RAG. Therefore, we select three representative dense retrieval models, Contriever, DPR and ANCE, as the target retrieval model (Gao and Callan, 2021; Karpukhin et al., 2020; Xiong et al., 2020). By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking.
(3) Manipulation target. For a controversial topic-queries-set $Q$, documents $doc_{tar}$ with the target opinion $S_{t}$ are manipulated by modifing $doc_{tar}$ through a series of subtle alterations $p_{adv}$, . This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3.
(4) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.
(5) Experimental parameters. In the knowledge-guided attack process, we set the maximum editing distance parameter $\epsilon$ to 0.2, the semantic similarity threshold $\lambda$ to 0.85, and the number of iterations $N$ to 5.For adversarial trigger generation, we configured the experimental parameters as follows: the number of beams was set to 3, the top-$k$ value to 10, the batch size to 32, the temperature value to 1.0, the learning rate to 0.005, and the sequence length to 10.
(6) Poisoned Target.For the PROCON dataset, to investigate the manipulation performance under more challenging conditions, we performed relevance ranking for the documents with respect to each topic-query set $Q$ and the target stance $S_{t}$ From the ranked list, we selected the last five documents (i.e., those with the lowest relevance) as the target poisoned documents. For the MS MARCO dataset, we utilized the top-1000 relevance-ranked passage list for each topic-query set. From this list, we selected the passage with the lowest average rank as the target passage. This approach ensures that the evaluation focuses on passages that are least relevant to the target queries, thus providing a more rigorous benchmark for the proposed method.
(7) Experimental environment. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
A.2. Details of baselines
PoisonedRAG: Zou et al. (Zou et al., 2024) propose PoisonedRAG which injects a few poisoned texts into the RAG database such that the LLM generates manipulated answers. They consider both black-box and white-box settings. In the black-box setting, they attempt to include the question itself as part of the poisoned text to make it semantically similar to the target question. For our task, we adapted this strategy to the topic-queries setting by randomly selecting one query from the topic-queries set $Q$ as the trigger and inserting it at the start of the document. The manipulated documents were then evaluated for their impact on retrieval rankings and subsequent comparisons.
PAT: As a representative adversarial retrieval attack strategy, PAT employs a pairwise generation paradigm. Given a target query, the target candidate item, and the top-ranked candidate item (referred to as the anchor, which guides adversarial text generation), the method utilizes gradient-based optimization of a pairwise loss function calculated between the target and the anchor. The strategy incorporates additional constraints for fluency and next-sentence prediction to ensure coherence. Using beam search, the final adversarial trigger, denoted as $T_{\text{pat}}$, is generated iteratively in an auto-regressive manner. In our task, we adapted PAT for topic-queries by generating $T_{\text{pat}}$ for the target document considering the whole topic-queries set and evaluating its effectiveness under black-box constraints.
Collision: It employs a gradient-based optimization approach to generate adversarial paragraphs, referred to as collisions, which are semantically similar to the target query. Using gradient descent, continuous representations of collisions are optimized and subsequently discretized into tokens via beam search. ASC provides three variants: aggressive, aggressive-regularized, and natural. For evaluation, we focused on the most and least aggressive variants to examine their effectiveness in generating adversarial texts. Using parameters and settings derived from the original Birch model implementation, collisions were generated with a fixed length of 10 tokens. In the context of topic-queries, ASC was applied to create collisions for the entire topic-queries set, and its transfer attack performance was evaluated on the black-box RAG retriever.
A.3. Ablation Study Supplement
Due to space limitations, only a subset of ablation study results is presented in the main text. Figure 10 illustrates the impact of the Polarity Control module in Topic-FlipRAG on the sentiment of the documents when the target polarity is set to PRO. Figures 12 and 13 depict the influence of various baseline methods on document sentiment.
Figure 10. Impact of Polarity Control on documents sentiment (target polarity:PRO) .
Figure 11. Impact of Polarity Control on document stance.
Figure 12. Impact of Polarity Control on document stance(target polarity:CON).
Figure 13. Impact of Polarity Control on document stance(target polarity:PRO).
Additionally, Figure 11 demonstrates the effect of the Polarity Control module on the actual stance polarity of the documents.The distinction between stance polarity and sentiment is critical. Stance polarity is determined by utilizing the opinion classification approach described in Section 9, which evaluates whether the document supports or opposes the corresponding topic. In contrast, sentiment is directly inferred using a sentiment classification model, specifically cardiffnlp/twitter-roberta-base-sentiment-latest.
 [^2] https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest
Table 8. Topic-Oriented RAG attack results on RPOCON dataset based on ANCE.
Method Target Stance ANCE+qwen-2.5-8B-instruct ANCE+llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
Clean CON 0.21 0.24 0.18 0.18 0.21 0.24 0.24 0.24 0.25 0.24
PRO 0.19 0.24 0.17 0.20 0.21 0.21 0.27 0.23 0.24 0.24
ASC CON 0.30 0.30 0.25 0.23 0.28 0.06 0.35 0.35 0.26 0.28 0.32 0.08
PRO 0.24 0.35 0.25 0.27 0.29 0.06 0.27 0.40 0.27 0.29 0.31 0.07
PAT CON 0.23 0.27 0.24 0.25 0.25 0.04 0.32 0.28 0.27 0.27 0.29 0.05
PRO 0.25 0.23 0.27 0.29 0.25 0.05 0.29 0.29 0.30 0.29 0.30 0.05
PoisonedRAG CON 0.26 0.34 0.36 0.44 0.35 0.13 0.32 0.43 0.36 0.40 0.36 0.09
PRO 0.31 0.37 0.30 0.28 0.33 0.12 0.34 0.31 0.30 0.31 0.33 0.09
Topic-FlipRAG CON 0.39 0.44 0.40 0.42 0.42 0.21 0.40 0.50 0.41 0.50 0.46 0.23
PRO 0.43 0.41 0.31 0.39 0.39 0.18 0.40 0.52 0.37 0.56 0.47 0.23
Table 9. Topic-Oriented RAG attack results on RPOCON dataset based on DPR.
Method Target Stance DPR+qwen-2.5-8B-instruct DPR+llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
Clean CON 0.22 0.26 0.27 0.25 0.25 0.25 0.24 0.21 0.23 0.23
PRO 0.23 0.26 0.24 0.25 0.24 0.24 0.27 0.22 0.22 0.24
Collision CON 0.25 0.30 0.28 0.21 0.27 0.02 0.33 0.34 0.33 0.31 0.33 0.10
PRO 0.27 0.34 0.31 0.29 0.31 0.07 0.28 0.39 0.39 0.39 0.35 0.12
PAT CON 0.22 0.23 0.34 0.26 0.26 0.01 0.32 0.26 0.35 0.19 0.28 0.05
PRO 0.23 0.23 0.23 0.23 0.25 0.01 0.21 0.35 0.35 0.27 0.29 0.04
PoisonedRAG CON 0.32 0.40 0.28 0.48 0.37 0.12 0.39 0.39 0.36 0.58 0.43 0.20
PRO 0.29 0.36 0.31 0.31 0.32 0.08 0.30 0.42 0.36 0.42 0.37 0.14
Topic-FlipRAG CON 0.45 0.42 0.40 0.49 0.44 0.19 0.45 0.48 0.41 0.48 0.46 0.23
PRO 0.41 0.55 0.40 0.39 0.46 0.21 0.37 0.52 0.42 0.42 0.44 0.21
A.4. RAG opinion manipulation result based on other retrievers
To further explore RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?, we conducted RAG-based opinion manipulation experiments on the PROCON dataset using ANCE and DPR as retrievers. The results obtained for each baseline are consistent with the overall trends observed in Table 3. Detailed experimental outcomes are provided in Table 8 and Table 9.